A firm has a cost of preferred stock 8%, and its yield to maturity is 10%. This firm has a tax rate of 35%. Given this information, which financing is cheaper for the firm: debt or preferred stock? A. Debt B. Preferred stock C. Same
Answer #1

Hi

Cost of preferred stock = 8%

Yield to maturity (cost of debt) is tax deductible.

after tax cost of debt = 10*(1-t) = 10*(1-35%)

= 6.5%

So Financing through debt is cheaper

Thanks